• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Yi Qiang

    Log entry time 08:15:50 on October 22, 2008

    Entry number 244330

    keyword=Target tests between Oct 21,6pm to Oct 22, 7am

    1. Target chamber pickup coil adjustment with water cell was done by 9:00 pm.
    2. At the meaning time, connected the heater power supplies and controller.
    3. Without any cell in the oven, the oven system was debuged and finally reached 235 C setting temperature, no sign of anything broken. Oven and the controller are working properly now. This was done by 4:00 am
    4. At the meaning time, the pumping chamber pickup coils are being adjusted, only the longitudinal pair was done. The transverse pair has fairly leak (probably due to the refection from Bigbite field clamp, so that one coil has to be titled quite a lot to cancel the residue RF signal)
    4. Locked the Hall for final laser alignment from 5:30 am
    5. Finished final alignment of vertical pumping, then stated the full laser power test, the test was interuppeted by a power trip of laser rack when the total power was increased to 100 W (20W for each laser).
    6. We decided to pause further test to let other people start their work. Therefore, at about 6:20 we shut down the heater and later one removed the 3He cell.

    A copy of this log entry has been emailed to: jpchen@jlab.org, yqiang@jlab.org, chiran@jlab.org, jkatich@jlab.org, jinhuang@jlab.org, zhangyi@jlab.org